OpenStudy (anonymous):

A hypothetical cube grows so that the length of its sides are increasing at a rate of 4 m/min. How fast is the volume of the cube increasing when the sides are 5m each?

OpenStudy (anonymous):

v=a^3 dv/dt=3a^2da/dt

OpenStudy (praxer):

let, side = a volume is \[a^3\] so dv/dt will give \[(3a^2)da/dt\]. since \[da/dt\] = 4m/min. now. dv/dt you can easily find right ??? @Nissa

OpenStudy (anonymous):

so it'll be (3x5m^2)(4m/min)= 300m/min? I'm sorry I haven't looked over this in a while.

OpenStudy (praxer):

no, it will be \[300m^3/min\]

OpenStudy (anonymous):

ohh okay, thank you

OpenStudy (praxer):

since it is the rate of change of volume...

OpenStudy (anonymous):

that makes sense now.